solve this system of equations by elimination
4x+y=-13
6x+-3y=-15

Answer by checkley77(12844)

4x+y=-13 MULTIPLY THIS EQUATION BY 3 & ADD THE 2 EQUATIONS.
6x-3y=-15 [I'LL ASSUME THE MIDDLE TERM IS A (-3)]
12X+3Y=-39
---------------
18X=-54
X=-54/18
X=-3 ANS.
4*-3+Y=-13
-12+Y=-13
Y=-13+12
Y=-1 ANS.
PROOF:
6*-3*-3*-1=-15
-18+3=15
15=15